OpenStudy (anonymous):

Write the answers as an improper fraction if possible. a.) 8 2/4-6 3/4= b.) 5 2/3-4 1/3= c.) 8 1/2+8 1/2= d.) 4 4/8+1 5/8=

OpenStudy (anonymous):

a) 34/4-27/4= 7/4 b) 17/3-13/3=4/3 c) 17/2+17/2= 34/2 d) 36/8+13/8= 49/8

OpenStudy (anonymous):

thank you doglover123